OpenStudy (ccchristian1):

Sue surveyed the students at her school to find out if they like sandwiches and/or tacos. The table below shows the results of the survey: Like Sandwiches Do Not Like Sandwiches Total Like Tacos 43 7 50 Do Not Like Tacos 27 23 50 Total 70 30 100 If a student likes sandwiches, what is the probability that student also likes tacos? 43% 61.4% 71.4% 86%

OpenStudy (ccchristian1):

@unkownbeast911

OpenStudy (ccchristian1):

if you don't understand that one Variable g is 5 more than variable w. Variable g is also 2 less than w. Which pair of equations best models the relationship between g and w? g = w + 5 g = w − 2 g = w − 5 g = w + 2 w = 5g w = g − 2 w = 5g w = g + 2

OpenStudy (anonymous):

I think it is B)

OpenStudy (anonymous):

i think they are both B

OpenStudy (ccchristian1):

for what question ?